OpenStudy (anonymous):

Find area enclosed by a given curves y=2x-x^2, y=2x-4

OpenStudy (anonymous):

first of all find their pt of intersection it is here (2,0) (-2,8) and then proceed

OpenStudy (anonymous):

how do you get that pt of intersection?

OpenStudy (antman):

set both equations equal to each other then solve for x. X=2 and x=-2.

OpenStudy (antman):

then use any equation and solve for y. (2,0) and (-2,-8)

OpenStudy (anonymous):

ic, then from here I will use integral from -8 to 0 with equation 2x-x^2 to find the area right?

OpenStudy (antman):

no, use \[\int\limits_{a}^{b} f(x)-g(x) dx \] if you use this form your limits are from [-2,2]

OpenStudy (antman):

f(x)= 2x-x^2 and g(x) =2x-4 a=-2 and b=2

OpenStudy (antman):

If you can graph this you can see that f(x) is above g(x).

OpenStudy (anonymous):

thanks ant man